### v2

# Bipartite entanglement extracted from multimode squeezed light generated in lossy waveguides

Abstract
Entangled two-mode Gaussian states constitute an important building block for continuous variable quantum computing and communication protocols and are thus of high demand for many experiments. In this work, we study such kind of states which are extracted from multimode light generated via type-II parametric down-conversion (PDC) in lossy waveguides. For such states, we demonstrate that the squeezing quantifies entanglement and we construct a measurement basis which results in the maximal bipartite entanglement. We illustrate our findings by numerically solving the spatial master equation for PDC in a Markovian environment. The optimal measurement modes are compared with two widely-used broadband bases: the Mercer-Wolf basis (the first-order coherence basis) and the Williamson-Euler basis.
Entangled two-mode bipartite Gaussian states play a fundamental role in quantum optics as they represent building blocks for many quantum optical protocols [1, 2]. An efficient source of such highly entangled states is thus an essential part for scalable quantum technological applications [3]. Currently, one of the most promising platforms for producing Gaussian quantum light is based on pulsed parametric down-conversion (PDC) in nonlinear waveguides [4, 5, 6, 7, 8]. However, two main effects are limiting the performance of such sources: scattering of guided light due to the surface roughness [9, 10] and the multimode structure of the generated PDC light [11, 12, 13, 14, 15]. Scattering losses are intrinsic waveguide losses that are caused by technological imperfections during the fabrication. Such type of losses results in the generation of a mixed state in the PDC process. The multimode PDC structure is induced by nonlinear optical coupling of interacting fields and strongly depends on the waveguide dispersion and the pump profile. Two-mode bipartite Gaussian states can be obtained from the initial multimode bipartite light by using a quantum pulse gate (QPG) [16, 17, 18] for each partition. Alternatively, one can selectively work with the chosen two modes of the multimode light by implementing proper gates with the use of homodyne detection [19, 20]. Both of these techniques require the determination of the correct mode shape, since measurements in the wrong modes can lead to the loss of non-classicality [21, 22], which becomes more crucial for PDC light generated in lossy waveguides.
In this paper, we study two-mode bipartite Gaussian states extracted from pulsed type-II multimode PDC generated in lossy waveguides. We show how to choose the measurement basis for the obtaining the maximally-entangled two-mode bipartite state (TMBS) which then can be used for a proper homodyne detection or QPG protocol.
Theoretical model.
For the description of PDC generated in lossy media we use the numerical scheme developed in Refs. [14, 15]. The details of this approach are given in the Appendices A and B. In the framework of multimode Gaussian states, the type-II PDC is described as a multimode bipartite system: part A (also called as signal) and part B (also called as idler), each consisting of $N$ bosonic monochromatic modes, $\hat{\mathbf{a}}=\big{(}\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big{)}^{T}$ and $\hat{\mathbf{b}}=\big{(}\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{N}\big{)}^{T}$, respectively. The lower indices correspond to the different frequencies. Having the vacuum state at the input of the waveguide, a solution of the spatial master equations the for correlation matrices for a type-II PDC process in a Markovian environment has the form [15], see also Appendix B,
$\mathcal{D}=\begin{pmatrix}\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}&\mathbf{0}\\ \mathbf{0}&\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}\end{pmatrix},~{}\mathcal{C}=\begin{pmatrix}\mathbf{0}&\braket{\mathbf{\hat{a}\hat{b}}}\\ \braket{\mathbf{\hat{b}\hat{a}}}&\mathbf{0}\end{pmatrix},$ (1)
where the matrices $\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}$, $\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}$ and $\braket{\mathbf{\hat{a}\hat{b}}}$ are the correlation matrices with elements $\braket{{\hat{a}^{\dagger}_{i}\hat{a}}_{j}}$, $\braket{{\hat{b}^{\dagger}_{i}\hat{b}_{j}}}$ and $\braket{{\hat{a}_{i}\hat{b}_{j}}}$, respectively. The expectation values $\braket{\mathbf{\hat{a}}^{\dagger}}=\braket{\mathbf{\hat{a}}}=\braket{\mathbf{\hat{b}}^{\dagger}}=\braket{\mathbf{\hat{b}}}\equiv 0$ due to the initial vacuum state. The relationship between the matrices $\mathcal{D}$, $\mathcal{C}$, and the covariance matrix $\Sigma$ are given in Appendix A.
Figure 1: The scheme of a TMBS preparation from a multimode type-II PDC process. After the lossy PDC process, a multimode mixed state with the covariance matrix $\Sigma$ is generated. Two quantum pulse gates (QPG) in partition A with the mode $\mathbf{u}_{A}$ and in partition B with the mode $\mathbf{u}_{B}$, respectively, provide the TMBS with the covariance matrix $\sigma(\mathbf{u}_{A},\mathbf{u}_{B})$.
To extract the maximally-entangled two-mode state from the initial multimode state, we have to find the proper shape of these two broadband modes, which will then be used in the QPGs as shown in Fig. 1 or in homodyne detection for realizing the desired continuous variable protocol. In the general case, such modes represent a superposition of the initial monochromatic PDC modes, therefore, mathematically they can be found using a passive local transformation $U=\big{(}\begin{smallmatrix}\mathbf{u}_{A}&\mathbf{0}_{N}\\ \mathbf{0}_{N}&\mathbf{u}_{B}\end{smallmatrix}\big{)}$, where $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ are normalized row-vectors of size $N$.
Then, the sought after broadband modes are $\hat{A}=\mathbf{u}_{A}\hat{\mathbf{a}}$ and $\hat{B}=\mathbf{u}_{B}\hat{\mathbf{b}}$. Applying the transformation $U$ to the initial multimode system Eq. (1), via the rules (10), one obtains the following nonzero second-order correlators for the introduced broadband modes:
$\braket{\hat{A}^{\dagger}\hat{A}}$ $=\mathbf{u}_{A}^{*}\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}\mathbf{u}_{A}^{T},$ (2)
$\braket{\hat{B}^{\dagger}\hat{B}}$ $=\mathbf{u}_{B}^{*}\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}\mathbf{u}_{B}^{T},$ (3)
$\braket{\hat{A}\hat{B}}$ $=\mathbf{u}_{A}\braket{\mathbf{\hat{a}\hat{b}}}\mathbf{u}_{B}^{T}.$ (4)
Note that the modes $\hat{A}$ and $\hat{B}$ can always be chosen to have $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$ via additional phase rotations $\mathbf{u}_{A}\rightarrow\mathbf{u}_{A}e^{i\phi_{A}}$ and $\mathbf{u}_{B}\rightarrow\mathbf{u}_{B}e^{i\phi_{B}}$. The number of photons in the signal and idler parts are $N_{A}=\braket{\hat{A}^{\dagger}\hat{A}}$ and $N_{B}=\braket{\hat{B}^{\dagger}\hat{B}}$, respectively.
By introducing the quadrature operators for broadband modes $\hat{q}^{A}=\hat{A}^{\dagger}+\hat{A}$, $\hat{p}^{A}=i(\hat{A}^{\dagger}-\hat{A})$ and $\hat{q}^{B}=\hat{B}^{\dagger}+\hat{B}$, $\hat{p}^{B}=i(\hat{B}^{\dagger}-\hat{B})$ (with the commutation relations $[\hat{q}^{A},\hat{p}^{A}]=[\hat{q}^{B},\hat{p}^{B}]=2i$, $[\hat{q}^{A},\hat{p}^{B}]=0$), one can construct the covariance matrix $\sigma(\mathbf{u}_{A},\mathbf{u}_{B})$. In what follows, we order the quadratures in this covariance matrix as $(\hat{q}^{A},\hat{p}^{A},\hat{q}^{B},\hat{p}^{B})$ leading to a matrix form, which is a special case of the so-called standard form [23, 24]
$\sigma(\mathbf{u}_{A},\mathbf{u}_{B})=\begin{pmatrix}\alpha&0&\gamma&0\\ 0&\alpha&0&-\gamma\\ \gamma&0&\beta&0\\ 0&-\gamma&0&\beta\\ \end{pmatrix},$ (5)
where $\alpha\equiv\braket{\hat{p}^{A}\hat{p}^{A}}=\braket{\hat{q}^{A}\hat{q}^{A}}=1+2\braket{\hat{A}^{\dagger}\hat{A}}$, $\beta\equiv\braket{\hat{p}^{B}\hat{p}^{B}}=\braket{\hat{q}^{B}\hat{q}^{B}}=1+2\braket{\hat{B}^{\dagger}\hat{B}}$ and $\gamma\equiv\braket{\hat{q}^{A}\hat{q}^{B}}=-\braket{\hat{p}^{A}\hat{p}^{B}}=2\braket{\hat{A}\hat{B}}$. Note that if $\textrm{Im}(\braket{\hat{A}\hat{B}})=0$, there are no correlations between the coordinates and momenta from different subsystems, namely $\braket{\hat{q}^{A}\hat{p}^{B}}=\braket{\hat{p}^{A}\hat{q}^{B}}=0$, see Appendix A.
To characterize the degree of entanglement of TMBS described by the matrix Eq. (5), we use the logarithmic negativity [25, 26] as the entanglement measure [27, 28]
$\mathcal{E}(\sigma)=\max\left[-\log\left(\nu_{-}(\tilde{\sigma})\right),0\right],$ (6)
where the value $\nu_{-}(\tilde{\sigma})$ is the minimal symplectic value of a partially transposed covariance matrix $\tilde{\sigma}=T\sigma T$ with $T=\mathrm{diag}(1,1)\oplus\mathrm{diag}(1,-1)$. Therefore, achieving the TMBS with the maximal degree of entanglement $E_{M}$ is reduced to maximizing the logarithmic negativity by varying the local transformation $U$:
$E_{M}=\max_{\begin{subarray}{c}\mathbf{u}_{A},\mathbf{u}_{B}:\\ |\mathbf{u}_{A}|=1,|\mathbf{u}_{B}|=1\end{subarray}}\mathcal{E}\left(\sigma(\mathbf{u}_{A},\mathbf{u}_{B})\right).$ (7)
The symplectic spectrum of the partially transformed covariance matrix ${\nu(\tilde{\sigma}})$ can be found as the eigenspectrum of the matrix $|i\Omega\tilde{\sigma}|$ [25, 27, 2], where $\Omega=\omega\oplus\omega$ and $\omega=\big{(}\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\big{)}$. The corresponding characteristic polynomial reads $p(x)=x^{4}-(\alpha^{2}+\beta^{2}+2\gamma^{2})x^{2}+(\alpha\beta-\gamma^{2})^{2}=0$ and results in the symplectic values
$\nu_{\pm}(\tilde{\sigma})=\frac{1}{2}\left(\alpha+\beta\pm\sqrt{(\alpha-\beta)^{2}+4\gamma^{2}}\right).$ (8)
On the other hand, the eigenvalues of the covariance matrix $\lambda(\sigma)$ are determined by the roots of its characteristic polynomial, which for the matrix Eq. (5) reads $p(\lambda)=\det(\sigma-\lambda\mathbf{1})=\left((\lambda-\alpha)(\lambda-\beta)-\gamma^{2}\right)^{2}=0$. As can be seen, its roots $\lambda_{\pm}(\sigma)$ are double degenerate and coincide with the symplectic spectrum of the partially transformed covariance matrix $\tilde{\sigma}$; namely, $\lambda_{\pm}(\sigma)\equiv\nu_{\pm}(\tilde{\sigma})$. At the same time, the smallest eigenvalue $\lambda_{-}(\sigma)$ determines the maximal degree of squeezing [29, 30] contained in the system. As long as for the studied system $\lambda_{-}(\sigma)=\nu_{-}(\tilde{\sigma})$, the maximal degree of squeezing is a proper entanglement measure.
This means that the optimization task Eq. (7) can be reformulated in terms of finding the maximal two-mode squeezing of the initial multimode state. Indeed, the lowest eigenvalue $\lambda_{min}$ of the covariance matrix $\Sigma$ determines the maximal squeezing present in the initial multimode system. Then, by determining the vectors $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$, for which $\lambda_{-}(\sigma(\mathbf{u}_{A},\mathbf{u}_{B}))=\lambda_{min}$, one can find the required shapes of the modes that maximize Eq. (7).
In general, the connection between squeezing and entanglement is not as trivial and was studied, e.g., in Ref. [30]: Squeezing is present in entangled Gaussian states, while the opposite is not always correct. However, in our system, there is no single-mode squeezing in either part A or B of the initial state. Indeed, the correlation matrices $\braket{\mathbf{\hat{a}\hat{a}}}=\braket{\mathbf{\hat{b}\hat{b}}}=\mathbf{0}$, see Eq. (1), result in the quadrature variances $\braket{(\Delta\hat{p}^{A})^{2}}\geq 1$ and $\braket{(\Delta\hat{p}^{B})^{2}}\geq 1$. These inequalities hold for any local transformation $U=\big{(}\begin{smallmatrix}\mathbf{u}_{A}&\mathbf{0}_{N}\\ \mathbf{0}_{N}&\mathbf{u}_{B}\end{smallmatrix}\big{)}$, therefore, the only possible squeezing $\braket{(\Delta\hat{p}^{F})^{2}}<1$ can be found only for some joint mode $\hat{F}=t\hat{A}+r\hat{B}$ with $|t|^{2}+|r|^{2}=1$ and $t,r\neq 0$. Such type of squeezing is known as two-mode squeezing. In turn, the mode $\hat{F}$ can always be chosen to get $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$.
Maximally-squeezed basis.
The basis that provides maximal squeezing of lossy systems and brings the covariance matrix to the second canonical form [29] is the maximally-squeezed basis (MSq-basis) [14]. The first mode $\hat{F}$ of this basis can be found from the eigenvector $\mathbf{v}_{min}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{N},y^{b}_{N})$ that corresponds to the eigenvalue $\lambda_{min}$, namely $\hat{F}=\sum_{n}\big{(}v^{a}_{n}\hat{a}_{n}+v^{b}_{n}\hat{b}_{n}\big{)}$, where we have introduced two complex vectors $[\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $[\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Then, the normalized vectors $\mathbf{u}^{MSq}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{MSq}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ give us the desired broadband modes that minimize Eq. (7). The vectors $\mathbf{u}^{MSq}_{A}$ and $\mathbf{u}^{MSq}_{B}$ define the state called MSq-TMBS and the corresponding covariance matrix $\sigma^{MSq}\equiv\sigma(\mathbf{u}^{MSq}_{A},\mathbf{u}^{MSq}_{B})$. The minimal eigenvalues of matrices $\Sigma$ and $\sigma^{MSq}$ coincide, which guarantees that the generated TMBS is a TMBS with maximal squeezing and, therefore, with maximal entanglement.
Studies of multimode properties of non-classical optical states are commonly performed with two other broadband bases: the Mercer-Wolf (MW) basis [31, 32] (also known as the first-order coherence mode basis [33]) and the Williamson-Euler (WE) basis [2, 34, 35, 36, 37, 38].
Based on the MW and WE decompositions, the vectors $\mathbf{u}^{MW}_{A}$, $\mathbf{u}^{MW}_{B}$ and $\mathbf{u}^{WE}_{A}$, $\mathbf{u}^{WE}_{B}$ can be obtained. These vectors define additional TMBSs (for details see Appendix C) that are further referred to as MW-TMBS and WE-TMBS with the covariance matrices $\sigma^{MW}\equiv\sigma(\mathbf{u}^{MW}_{A},\mathbf{u}^{MW}_{B})$ and $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$, respectively.
The special case where the Mercer-Wolf, Williamson-Euler, and MSq bases coincide corresponds to a pure multimode state, i.e., to PDC generated in a lossless waveguide [14]. This basis is known as the Schmidt-mode basis [33, 39]. The first Schmidt mode is a mode with the highest number of photons and, at the same time, with the smallest eigenvalue of the covariance matrix. Building the TMBS based on the first Schmidt mode, we obtain a pure state that is characterized by a covariance matrix $\sigma^{S}$ with parameters $\alpha=\beta$ and $\gamma=\sqrt{1-\alpha^{2}}$ and corresponds to the ideal two-mode squeezer.
Numerical simulations.
Figure 2: Numerical results for a lossless waveguide, the parameters of which are given in the main text. (a) Normalized joint spectral intensity for low-gain PDC, $\delta\omega=\omega-\omega_{0}$, where $\omega_{0}=\omega_{p}/2$ is the central frequency of the PDC light. (b) Dependence of the smallest quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$ as a function of the number of photons $N_{A}=N_{B}$. The additional axis shows the corresponding logarithmic negativity $\mathcal{E}(\sigma^{S})$.
Figure 3: Numerical results for lossy waveguides: (a,c,e) WG1 and (b,d,f) WG2. The waveguide parameters are given in the main text. Different colors correspond to TMBS build with different modes: MW (blue), WE (green), and MSq (red). (a, b) The number of photons $N_{A}$ (solid lines) and $N_{B}$ (dashed lines) as a function of losses $\bar{\eta}$. (c, d) The dependencies of the smallest quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}=\lambda_{-}(\sigma)$ for $\sigma^{MW}$, $\sigma^{WE}$ and $\sigma^{MSq}$ as a function of losses $\bar{\eta}$. Additional right axis show the logarithmic negativity $\mathcal{E}(\sigma)$. The gray area highlights the region with zero logarithmic negativity. (e, f) Purities of $\sigma^{MW}$, $\sigma^{WE}$ and $\sigma^{MSq}$ as a function of losses $\bar{\eta}$.
To simulate frequency-degenerate type-II PDC, we consider a $1$ cm long waveguide with manually defined dispersion and losses. As the pump, we use a Gaussian pulse with a full width at half maximum of $\Delta\tau=0.4$ ps and a central wavelength of $\lambda_{p}=755$ nm. For simplicity, we limit ourselves to the linear refractive index dependence for each field, which is parametrized as $n(\omega)=n(\omega_{0})+\frac{\omega-\omega_{0}}{\omega_{0}}\left[\frac{c}{v_{g}(\omega_{0})}-n(\omega_{0})\right],$ where $c$ is the speed of light, $v_{g}$ is the group velocity, and $\omega_{0}$ is the central frequency. For our studies, we choose the following parameters for the pump, signal, and idler fields (indices $p$, $s$ and $i$, respectively): $n_{p}=n(\omega_{p})=1.9$, $v_{g}^{p}=0.9c/n_{p}$; $n_{s}=n(\omega_{p}/2)=1.9$, $v_{g}^{s}=0.96v_{g}^{p}$; $n_{i}=n(\omega_{p}/2)=1.8$ $v_{g}^{i}=0.98v_{g}^{p}$. Quasi-phase-matching is obtained with $k_{QPM}=\frac{\omega_{p}}{2c}(2n_{p}-n_{s}-n_{i})$.
Usually, the internal waveguide losses for the TE- and the TM-modes are different [10]. In order to take this into account, we introduce two frequency-independent loss coefficients for the signal and idler fields $\eta_{s}$ and $\eta_{i}$, respectively. These coefficients can be parametrized as $\bar{\eta}=\frac{\eta_{s}+\eta_{i}}{2},\,r_{\eta}=\frac{\eta_{s}-\eta_{i}}{\eta_{s}+\eta_{i}}.$ In this paper, we consider three waveguides:
  • lossless waveguide WG0 with $\eta_{s}=\eta_{i}=0$;
  • waveguide WG1 with equal losses $\eta_{s}=\eta_{i}$, $r_{\eta}=0$;
  • waveguide WG2 with unbalanced losses for the signal and idler modes $\eta_{s}=2\eta_{i}$, $r_{\eta}=\frac{1}{3}$.
In the limit $\bar{\eta}\rightarrow 0$, both waveguides WG1 and WG2 coincide with the lossless waveguide WG0. We numerically solve the master equations (see Appendix B) for each waveguide for different parametric gains $\Gamma$ and losses $\bar{\eta}$, $r_{\eta}$. For the resulting correlation matrices $\mathcal{D}(\Gamma,\bar{\eta},r_{\eta})$ and $\mathcal{C}(\Gamma,\bar{\eta},r_{\eta})$, we find the Mercer-Wolf, Williamson-Euler, and MSq modes. For each basis, we calculate the TMBS with the algorithms described above.
Figure 4: TMBS for the lossy waveguide WG2 (unbalanced losses) with $\bar{\eta}=5$ dB and $r=\frac{1}{3}$. (a,b,c) Covariance matrices $\sigma^{MW}$, $\sigma^{WE}$ and $\sigma^{MSq}$, respectively. (d) Absolute value and (e) phase of signal modes $\mathbf{u}_{A}$; (f) absolute value and (g) phase of idler modes $\mathbf{u}_{B}$. The blue, green, and red colors correspond to the MW-, WE-, and MSq-TMBS bases, respectively.
Fig. 2 shows results of numerical simulations for the lossless waveguide WG0, for which, as was mentioned above, there exists the optimal TMBS based on the first Schmidt mode. Fig. 2(b) demonstrates the dependence of the smallest eigenvalue $\lambda_{-}(\sigma^{S})$ (characterized by quadrature variance $\braket{(\Delta\hat{p}^{F})^{2}}$) as a function of the number of photons. The obtained TMBS is pure and corresponds to the two-mode squeezed state; therefore, entanglement and squeezing are always present in this state.
In Fig. 3, results of numerical simulations for the lossy waveguides WG1 and WG2 are presented. The simulations were performed for a fixed gain $\Gamma$, which for the corresponding lossless waveguide WG0 gives the number of photons $N_{A}=N_{B}=40$ and the logarithmic negativity $\mathcal{E}(\sigma^{S})=5.1$. In contrast to the lossless case, here the MW-, WE-, and MSq-TMBS provide different quantum states. In Fig. 3(a, b) the number of photons for the signal and idler parts for different TMBSs are calculated as a function of internal losses $\bar{\eta}$. As can be seen, with increasing $\bar{\eta}$ the number of photons in each TMBS decreases. For the waveguides with equal losses, the number of photons in the signal and idler subsystems coincides, while the unbalanced losses lead to different numbers of signal and idler photons in each TMBS. Note that the MW-TMBS contains the maximally possible number of photons in each partition. Figs. 3(c) and (d) demonstrate the dependence of the lowest eigenvalues $\lambda_{-}(\sigma^{MW})$, $\lambda_{-}(\sigma^{WE})$, and $\lambda_{-}(\sigma^{MSq})$ on the internal loss coefficient $\bar{\eta}$ for WG1 and WG2, respectively. As expected, the increase of the internal losses $\bar{\eta}$ decreases both the squeezing and entanglement of all studied TMBS. However, the maximally entangled state is clearly realized for the MSq-TMBS. Moreover, the purity of MSq-TMBS is the highest compared to $\sigma^{MW}$ and $\sigma^{WE}$ as is demonstrated in Fig. 3(e,f).
In the case of unbalanced losses, choosing a wrong basis may even lead to a loss of entanglement, as it is demonstrated in Fig. 3(d) for WG2. Here, for $\bar{\eta}>2.6$ dB, the lowest eigenvalue is grater than unity, $\lambda_{-}(\sigma^{MW})>1$, which means that MW-TMBS becomes separable (see the gray region in Fig. 3(d)). For the losses that exhibit the maximum difference in the smallest eigenvalue between the three bases in Fig. 3(d), namely, $\bar{\eta}=5$ dB, we present the explicit form of the covariance matrices $\sigma^{MW}$, $\sigma^{WE}$, and $\sigma^{MSq}$, see Fig. 4(a-c), as well as the spectral and phase profiles of $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ modes, see Fig. 4(d-g). It is clearly seen that the structure of the covariance matrix and the spectral width of the modes strongly depend on the chosen basis. For example, in the MW-basis, the modes $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ have almost the same width, while in the MSq-basis their widths differ significantly. The mean number of photons, the lowest eigenvalue, as well as the negativity and purity for this special loss point are presented in Table 1 and show the importance of choosing the correct basis for a particular lossy system.
Summing up, we demonstrate how the TMBS with the maximal degree of squeezing can be created from a pulsed type-II multimode PDC process in lossy waveguides. For the obtained TMBS, we show that the maximal degree of squeezing quantifies entanglement and we introduce a basis (MSq-basis) in which the TMBS with the maximal entanglement can be constructed. We compare this optimal TMBS with the TMBS obtained via the Mercer-Wolf and the Williamson-Euler modes. By numerical simulations of lossy waveguides, we demonstrate that neither the first Mercer-Wolf mode, nor the first Williamson-Euler mode are optimal for preparing the maximally entangled state. Therefore, an intuitive experimental approach of maximizing intensity would not result in an optimal entangled TMBS.
From an experimental point of view, to find the MSq-basis, one can either implement a multiparameter optimization procedure via Eq. (7) using the two-mode squeezing $\lambda_{-}\left(\sigma(\mathbf{u}_{A},\mathbf{u}_{B})\right)$ as a target parameter or provide a diagonalization of the measured covariance matrix in order to find the proper shapes of $\mathbf{u}_{A}$ and $\mathbf{u}_{B}$ modes. Herewith, a complete classical waveguide characterization can significantly simplify the multiparameter optimization task by limiting the range of parameters.
TMBS Mer.-Wolf Will.-Euler MSq
$N_{A}$ 20.6 16.3 10.2
$N_{B}$ 24.7 17.9 10.8
$\lambda_{-}(\sigma)=\nu_{-}(\tilde{\sigma})$ 1.55 0.12 0.10
$\braket{(\Delta\hat{p})^{2}}$, dB 1.90 -9.07 -9.90
$\mathcal{E}(\sigma)$ 0 2.09 2.28
Purity 0.007 0.115 0.223
Table 1: The results for different TMBS for the lossy waveguide WG2 (unbalanced losses) with $\bar{\eta}=5$ dB and $r=\frac{1}{3}$.
Acknowledgements.
This work is supported by the ‘Photonic Quantum Computing’ (PhoQC) project, which is funded by the Ministry for Culture and Science of the State of North-Rhine Westphalia. We also acknowledge financial support of the Deutsche Forschungsgemeinschaft (DFG) via the TRR 142/3 (Project No. 231447078, Subproject No. C10). We acknowledge support for the publication cost by the Open Access Publication Fund of Paderborn University.
Appendix A Multimode bipartite Gaussian state
A detailed descriptions can be found in Refs. [14, 15]. Let us consider the multimode Gaussian system, consisting of two parts: the part A, which is consists of $N$ bosonic modes, denoted as $\hat{\mathbf{a}}=\big{(}\hat{a}_{0},\hat{a}_{1},\dots,\hat{a}_{N}\big{)}^{T}$ and the part B, consisting of $M$ bosonic modes $\hat{\mathbf{b}}=\big{(}\hat{b}_{0},\hat{b}_{1},\dots,\hat{b}_{M}\big{)}^{T}$. Commutation relations are bosonic $[\hat{a}_{i}(z),\hat{a}^{\dagger}_{j}(z)]=[\hat{b}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=\delta_{ij}$ and $[\hat{a}_{i}(z),\hat{b}^{\dagger}_{j}(z)]=0$. The correlation matrices that fully describe the state of the full system read
$\mathcal{D}_{ab}=\begin{pmatrix}\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}&\braket{\mathbf{\hat{a}^{\dagger}\hat{b}}}\\ \braket{\mathbf{\hat{b}^{\dagger}\hat{a}}}&\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}\end{pmatrix},~{}\mathcal{C}_{ab}=\begin{pmatrix}\braket{\mathbf{\hat{a}\hat{a}}}&\braket{\mathbf{\hat{a}\hat{b}}}\\ \braket{\mathbf{\hat{b}\hat{a}}}&\braket{\mathbf{\hat{b}\hat{b}}}\end{pmatrix}.$ (9)
The expressions in the form $\braket{\mathbf{\hat{a}^{\dagger}\hat{b}}}$ and $\braket{\mathbf{\hat{a}\hat{b}}}$ denote the $N\times M$ matrices with matrix elements $\braket{\hat{a}_{i}^{\dagger}(z)\hat{b}_{j}(z)}$ and $\braket{\hat{a}_{i}(z)\hat{b}_{j}(z)}$, respectively.
The general local passive unitary transformation is defined as $U=U_{A}\oplus U_{B}$, where the new broadband modes are $\hat{\mathbf{A}}=U_{A}\hat{\mathbf{a}}$ and $\hat{\mathbf{B}}=U_{B}\hat{\mathbf{b}}$. The correlation matrices in the new broadband modes are
$\mathcal{D}_{AB}=U^{*}\mathcal{D}_{ab}U^{T},\text{ ~{} and ~{} }\mathcal{C}_{AB}=U\mathcal{C}_{ab}U^{T}.$ (10)
In addition to the creation and annihilation operators, we also need the quadrature operators. For arbitrary modes with the annihilation operator $\hat{F}$, the quadrature operators are $\hat{q}^{F}_{i}=\hat{F}^{\dagger}_{i}+\hat{F}_{i}$ and $\hat{p}^{F}_{i}=i(\hat{F}^{\dagger}_{i}-\hat{F}_{i})$ with the commutation relations $[\hat{q}^{F}_{n},\hat{p}^{F}_{m}]=2i\delta_{nm}$.
The quadrature operators allow us to introduce the covariance matrix for the full system.
$\sigma_{ij}=\dfrac{\braket{\hat{x}_{i}\hat{x}_{j}+\hat{x}_{j}\hat{x}_{i}}}{2}-\braket{\hat{x}_{i}}\braket{\hat{x}_{j}},$ (11)
with the vector $\hat{\mathbf{x}}=(\hat{q}^{a}_{1},\hat{p}^{a}_{1},\dots,\hat{q}^{a}_{N},\hat{p}^{a}_{N};\hat{q}^{b}_{1},\hat{p}^{b}_{1},\dots,\hat{q}^{b}_{M},\hat{p}^{b}_{M})^{T}$, while the quadrature operators are $\hat{q}^{c}_{i}=\hat{c}^{\dagger}_{i}+\hat{c}_{i}$ and $\hat{p}^{c}_{i}=i(\hat{c}^{\dagger}_{i}-\hat{c}_{i})$ with the commutation relations $[\hat{q}^{c}_{n},\hat{p}^{d}_{m}]=2i\delta_{nm}\delta_{cd}$. Here, the upper indexes indicate arbitrary modes $c,d=(a,b)$. The elements of the covariance matrix are connected with the correlation matrices (9) as
$\braket{\hat{q}^{c}_{i}\hat{q}^{d}_{j}}=\delta_{ij}\delta_{cd}+2\Big{(}\text{Re}[\braket{\hat{c}^{\dagger}_{i}\hat{d}_{j}}]+\text{Re}[\braket{\hat{c}_{i}\hat{d}_{j}}]\Big{)},$ (12)
$\braket{\hat{p}^{c}_{i}\hat{p}^{d}_{j}}=\delta_{ij}\delta_{cd}+2\Big{(}\text{Re}[\braket{\hat{c}^{\dagger}_{i}\hat{d}_{j}}]-\text{Re}[\braket{\hat{c}_{i}\hat{d}_{j}}]\Big{)},$ (13)
$\dfrac{\braket{\hat{p}^{c}_{i}\hat{q}^{d}_{j}+\hat{q}^{d}_{j}\hat{p}^{c}_{i}}}{2}=2\Big{(}\text{Im}[\braket{\hat{c}_{i}\hat{d}_{j}}]-\text{Im}[\braket{\hat{c}^{\dagger}_{i}\hat{d}_{j}}]\Big{)},$ (14)
$\braket{\hat{q}^{c}_{i}}=2\text{Re}[\braket{\hat{c}_{i}}],~{}~{}~{}\braket{\hat{p}^{c}_{j}}=2\text{Im}[\braket{\hat{c}_{j}}].$ (15)
Appendix B Master equation
The type-II PDC in a lossy waveguide can be described in terms of a master equation of the correlation matrices (9). The signal and idler fields are taken in a discrete uniform frequency space $(\omega_{0},\omega_{1},\dots,\omega_{N})$, which allows us to describe signal and idler systems by $N$ monochromatic modes $\hat{\mathbf{a}}(z)$ and $N$ monochromatic modes $\hat{\mathbf{b}}(z)$, respectively. The spatial master equation for PDC correlation matrices has the form
$\dfrac{d\mathcal{D}(z)}{dz}=i\big{(}\mathcal{D}(z)K-K^{*}\mathcal{D}(z)\big{)}\\ +i\Gamma\big{(}\mathcal{C}^{*}(z)M^{T}(z)-M^{*}(z)\mathcal{C}(z)\big{)},$ (16)
$\dfrac{d\mathcal{C}(z)}{dz}=i\big{(}\mathcal{C}(z)K+K\mathcal{C}(z)\big{)}\\ +i\Gamma\Big{(}\big{(}M(z)\mathcal{D}(z)+M(z)\big{)}^{T}+M(z)\mathcal{D}(z)\Big{)},$ (17)
where the superscript $[\cdot]^{*}$ denotes the complex conjugation of a matrix. The matrix $K$ is a diagonal matrix consisting of complex wave vectors $\textrm{diag}(\kappa^{a}_{0},\dots\kappa^{a}_{N},\kappa^{b}_{0},\dots,\kappa^{b}_{N})$. Here, $\kappa^{a}_{n}=k^{a}_{n}+i\eta_{a}/2$, where $k^{a}_{n}$ is a real wave vector for mode $\hat{a}_{n}$ and $\eta_{a}$ is the loss coefficient for subsystem $a$. The $z$-dependent matrix $M(z)$ is given by
$M(z)=\begin{pmatrix}\mathbf{0}_{N}&J(z)\\ J^{T}(z)&\mathbf{0}_{N}\end{pmatrix}$ (18)
and is a coupling matrix of a spatial generator for the PDC process $\hat{G}_{pdc}(z)=\frac{1}{2}\hbar\Gamma\sum_{i,j}J_{ij}(z)\hat{a}_{i}^{\dagger}(z)\hat{b}_{j}^{\dagger}(z)+h.c.$ Explicitly, the matrix $J_{ij}(z)=S(\omega_{i}+\omega_{j})e^{i(k_{p}(\omega_{i}+\omega_{j})-k_{QPM})z}$; $S(\omega)$ is the pump spectral distribution at $z=0$; $k^{(a,b,p)}_{n}\equiv k^{(a,b,p)}(\omega_{n})=\frac{n^{(a,b,p)}(\omega_{n})\omega_{n}}{c}$ are the wave vectors of the (a) signal, (b) idler, and (p) pump fields in the waveguide; $k_{QPM}=2\pi/\Lambda$ and $\Lambda$ is the poling period for the quasi-phase-matching condition. The parameter $\Gamma$ determines the coupling strength of the PDC process.
The initial condition (vacuum state) reads $\mathcal{D}(0)=\mathcal{C}(0)=\mathbf{0}_{2N}$ and, together with the coupling matrix in the form of Eq. (18), determines the structure of the solution in the form of Eq. (1).
Appendix C Mercer-Wolf and Williamson-Euler decompositions
Mercer-Wolf modes
can be obtained via the diagonalization of the matrix $\mathcal{D}$ [14, 15]. Its diagonalization can be obtained via the independent eigendecomposition of the matrices $\braket{\mathbf{\hat{a}^{\dagger}\hat{a}}}=V_{A}\Lambda_{A}V_{A}^{H}$ and $\braket{\mathbf{\hat{b}^{\dagger}\hat{b}}}=V_{B}\Lambda_{B}V_{B}^{H}$, where the diagonal matrices $\Lambda_{A}$ and $\Lambda_{B}$ are given in decreasing order.
The TMBS, based on the Mercer-Wolf modes (MW-TMBS) is built for the vectors $\mathbf{u}^{MW}_{A}=e^{i\frac{\phi}{2}}\mathbf{v}_{A}$ and $\mathbf{u}^{MW}_{B}=e^{i\frac{\phi}{2}}\mathbf{v}_{B}$, where the row-vectors $\mathbf{v}_{A}$ and $\mathbf{v}_{B}$ are the first rows of the matrices $V_{A}^{T}$ and $V_{B}^{T}$, respectively, and a constant phase $\phi=\mathrm{arg}\big{(}\mathbf{v}_{A}\braket{\mathbf{\hat{a}\hat{b}}}\mathbf{v}_{B}^{T}\big{)}$. The covariance matrix of MW-TMBS is defined as $\sigma^{MW}\equiv\sigma(\mathbf{u}^{MW}_{A},\mathbf{u}^{MW}_{B})$. As long as the modes are obtained via the diagonalization of matrix $\mathcal{D}$, the first Mercer-Wolf modes $\mathbf{u}^{MW}_{A}$ and $\mathbf{u}^{MW}_{B}$ contain the maximally possible number of photons per mode.
Williamson-Euler modes
are obtained via the Williamson-Euler decomposition [37, 14], namely a decomposition of the initial covariance matrix $\sigma=O_{l}\Lambda O_{r}DO_{r}^{T}\Lambda O_{l}^{T}$, where the matrices $O_{l}$ and $O_{r}$ are orthogonal and matrices $D=\text{diag}(\nu_{1},\nu_{1},\dots,\nu_{N},\nu_{N})$ and $\Lambda=\text{diag}(e^{r_{1}},e^{-r_{1}},\dots,e^{r_{N}},e^{-r_{N}})$ are diagonal; the values $r_{i}$ are sorted in descending order. Having the first vector $\mathbf{v}^{(1)}=(x^{a}_{1},y^{a}_{1},\dots,x^{a}_{N},y^{a}_{N};x^{b}_{1},y^{b}_{1},\dots,x^{b}_{M},y^{b}_{M})$ of matrix $O_{l}$, and applying the same procedure as for the MSq mode, we can build modes for each part as $\hat{A}^{WE}=\mathbf{u}^{WE}_{A}\mathbf{\hat{a}}$ and $\hat{B}^{WE}=\mathbf{u}^{WE}_{B}\mathbf{\hat{b}}$, where $\mathbf{u}^{WE}_{A}=\frac{\mathbf{v}^{a}}{|\mathbf{v}^{a}|}$ and $\mathbf{u}^{WE}_{B}=\frac{\mathbf{v}^{b}}{|\mathbf{v}^{b}|}$ with $[\mathbf{v}^{a}]_{n}=y^{a}_{n}+ix^{a}_{n}$ and $[\mathbf{v}^{b}]_{n}=y^{b}_{n}+ix^{b}_{n}$. Numerically, the Williamson-Euler decomposition can be found with the use of algorithms presented in, e.g., Refs. [35, 36, 38]. The TMBS obtained by the vectors $\mathbf{u}^{WE}_{A}$ and $\mathbf{u}^{WE}_{B}$ is further called as WE-TMBS; the corresponding covariance matrix is defined as $\sigma^{WE}\equiv\sigma(\mathbf{u}^{WE}_{A},\mathbf{u}^{WE}_{B})$.